# The FFT Strikes Back: An Efficient Alternative to Self-Attention

### v3

@@ Abstract
Conventional self-attention mechanisms incur quadratic complexity, limiting their scalability on long sequences. We introduce FFTNet, an adaptive spectral filtering framework that leverages the Fast Fourier Transform (FFT) to achieve global token mixing in $\mathcal{O}(n\log n)$ time. By transforming inputs into the frequency domain, FFTNet exploits the orthogonality and energy preservation guaranteed by Parseval’s theorem to capture long-range dependencies efficiently. A learnable spectral filter and modReLU activation dynamically emphasize salient frequency components, providing a rigorous and adaptive alternative to traditional self-attention. Experiments on the Long Range Arena and ImageNet benchmarks validate our theoretical insights and demonstrate superior performance over fixed Fourier and standard attention models.
@@ 3 Adaptive Spectral Filtering Method
In this section, we present an adaptive spectral filtering framework that eliminates the need for explicit pairwise interactions in global token mixing. Instead of relying on dot-product-based self-attention, we employ the discrete Fourier transform (DFT) to capture long-range dependencies efficiently. By adaptively modulating the resulting frequency components and then applying an inverse transform, our method strikes an effective balance between expressive power and computational cost, supported by strong theoretical guarantees.
@@ 3.1 Motivation
Standard attention mechanisms compute pairwise interactions between tokens, incurring a quadratic cost in sequence length. As the number of tokens grows, this approach rapidly becomes prohibitive. In contrast, the Fourier transform decomposes a sequence into frequency components, inherently encoding global interactions in $\mathcal{O}(n\log n)$ time. To enhance representational power, we introduce a learnable filter in the frequency domain, enabling the model to emphasize salient frequency bands while retaining computational efficiency.
@@ 3.2 Method Description
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be the input sequence of length $n$ and embedding dimension $d$. Our method comprises four steps:
1. Fourier Transform.
We first apply the discrete Fourier transform along the token dimension:
$\mathbf{F}=\operatorname{FFT}(\mathbf{X})\;\in\;\mathbb{C}^{n\times d}.$
This operation represents each embedding across orthogonal frequency components, enabling global interactions without explicit pairwise comparisons.
@@
To selectively emphasize important frequencies, we use a learnable filter. First, compute a global context vector:
$\mathbf{c}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i}.$
This context is fed into a multi-layer perceptron that outputs two sets of modulation parameters defined over the frequency bins and attention heads: a multiplicative adaptive scale $\Delta\mathbf{s}$ and an additive adaptive bias $\Delta\mathbf{b}$, with $\Delta\mathbf{s},\,\Delta\mathbf{b}\in\mathbb{R}^{H\times f\times 1}$ (where $H$ is the number of heads and $f=\lfloor\tfrac{n}{2}\rfloor+1$ denotes the frequency bins). With a fixed base filter $\mathbf{W}_{\mathrm{base}}$ and a base bias $\mathbf{b}_{\mathrm{base}}$, the effective parameters are defined as:
$\mathbf{W}=\mathbf{W}_{\mathrm{base}}\odot\bigl{(}1+\Delta\mathbf{s}\bigr{)},\qquad\mathbf{b}=\mathbf{b}_{\mathrm{base}}+\Delta\mathbf{b}.$
The adaptive filtering step is then performed in the frequency domain:
$\tilde{\mathbf{F}}=\mathbf{F}\odot\mathbf{W}+\mathbf{b},$
which reweights and shifts the Fourier coefficients according to the global context.
3. Nonlinear Activation (modReLU).
To capture higher-order relationships in the complex frequency domain, we apply the modReLU Arjovsky et al. (2016) activation, defined for a complex number $z=re^{i\theta}$ (with $r=|z|$ and $\theta=\arg(z)$) as:
$\operatorname{modReLU}(z)\;=\;\begin{cases}(r+b)\,e^{i\theta},&\text{if }r+b>0,\\[6.0pt] 0,&\text{otherwise},\end{cases}$
where $b$ is a learnable bias. This operation applies a ReLU-like threshold to the magnitude while preserving the phase, making it well suited for frequency-domain representations. The element-wise modReLU is then:
$\tilde{\mathbf{F}}=\operatorname{modReLU}\bigl{(}\tilde{\mathbf{F}}\bigr{)}.$
4. Inverse Fourier Transform.
Finally, we return to the token domain via the inverse Fourier transform, retaining the real component of the reconstructed signal:
@@
$\mathbf{Y}=\operatorname{IFFT}\bigl{(}\tilde{\mathbf{F}}\bigr{)}\;\in\;\mathbb{R}^{n\times d}.$
This yields a globally mixed representation that incorporates adaptive filtering and nonlinear transformations in the frequency domain.
3.3 Theoretical Justification for FFT over Self-Attention
Our FFT-based adaptive spectral filtering approach offers several advantages over conventional self-attention:
  • Efficient Global Mixing: By decomposing inputs into frequency components, the FFT provides global interactions in $\mathcal{O}(n\log n)$ time—much more scalable than the $\mathcal{O}(n^{2})$ of self-attention.
  • Implicit, Adaptive Attention: The adaptive spectral filter effectively learns a frequency-domain mask informed by the global context vector. This mask reweights crucial frequency bands similarly to attention weights but avoids explicit pairwise computations.
  • Greater Expressivity via Nonlinearity: While the Fourier transform itself is linear, applying modReLU to the complex coefficients enriches representational capacity. This allows the model to capture intricate, higher-order patterns that purely linear operations may miss.
  • Energy Preservation and Stability: Parseval’s theorem ensures the input signal’s norm (energy) is preserved by the FFT. Thus, our method maintains stability by avoiding accidental loss of crucial information.
By leveraging frequency-domain global mixing, learnable spectral filtering, and the modReLU nonlinear activation, our method is both theoretically grounded and computationally efficient, serving as a robust alternative to self-attention.
3.4 Computational Complexity
The central operations in our method are the Fast Fourier Transform (FFT) and its inverse, each costing $\mathcal{O}(n\log n)$ per channel. In contrast, self-attention requires $\mathcal{O}(n^{2})$ time for pairwise computations. The adaptive filtering and activation steps introduce only a linear overhead $\mathcal{O}(n)$, keeping the overall complexity at $\mathcal{O}(n\log n)$.
3.5 Summary
In essence, adaptive spectral filtering transforms the input into the frequency domain, applies a learnable nonlinear modulation (modReLU), and then inverts the transform to obtain a globally mixed representation. This provides:
  • A favorable $\mathcal{O}(n\log n)$ complexity.
  • Robust mechanisms for modeling long-range dependencies.
  • An efficient, expressive alternative to self-attention.
The code is available here: https://github.com/jacobfa/fft.
@@ 3.6 Proof of Computational Complexity
Fourier and Inverse Fourier Transforms.
For $\mathbf{X}\in\mathbb{R}^{n\times d}$, computing the FFT and IFFT requires $\mathcal{O}(n\log n)$ operations per channel, totaling $\mathcal{O}(d\cdot n\log n)$.
Adaptive Spectral Filtering.
  • Summation over $n$ tokens for the context vector $\mathbf{c}$ takes $\mathcal{O}(n)$.
  • Computing the adaptive modulation parameters via the MLP introduces negligible overhead.
  • Element-wise filtering on $\mathbf{F}$ is $\mathcal{O}(n)$ per channel.
Overall, filtering costs $\mathcal{O}(d\cdot n)$.
Nonlinear Activation.
Applying modReLU on all complex elements is $\mathcal{O}(n)$ per channel, or $\mathcal{O}(d\cdot n)$ in total.
@@
Since $\mathcal{O}(d\cdot n\log n)$ (for FFT/IFFT) dominates $\mathcal{O}(d\cdot n)$ (for filtering and activation), the total cost is $\mathcal{O}(d\cdot n\log n)$. For most practical settings where $d$ is not far larger than $n$, this effectively behaves as $\mathcal{O}(n\log n)$.
3.7 Proofs and Theoretical Guarantees
Below, we show key theoretical properties that justify our method as an efficient surrogate for self-attention.
3.7.1 Global Mixing via Orthogonal Decomposition
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be the input and $\mathbf{F}=\operatorname{FFT}(\mathbf{X})$. A unitary DFT matrix $\mathbf{F}_{n}\in\mathbb{C}^{n\times n}$ satisfies
$\mathbf{F}_{n}^{*}\mathbf{F}_{n}=n\mathbf{I},$
where $\mathbf{F}_{n}^{*}$ is the conjugate transpose and $\mathbf{I}$ is the identity. This orthogonality preserves inner products:
$\langle\mathbf{X}_{i},\mathbf{X}_{j}\rangle\;=\;\frac{1}{n}\langle\mathbf{F}_{i},\mathbf{F}_{j}\rangle,$
implying that each token influences every frequency component. Consequently, the transformation encodes global interactions much like self-attention, without $\mathcal{O}(n^{2})$ pairwise computations.
3.7.2 Energy Preservation via Parseval’s Theorem
Parseval’s theorem states:
$\|\mathbf{X}\|_{2}^{2}\;=\;\frac{1}{n}\,\|\mathbf{F}\|_{2}^{2}.$
After adaptive filtering and activation, the output is
$\mathbf{Y}=\operatorname{IFFT}\Bigl{(}\operatorname{modReLU}\bigl{(}\mathbf{F}\odot\mathbf{W}\bigr{)}\Bigr{)}.$
Provided the filter does not excessively amplify particular frequencies, the energy of $\mathbf{Y}$ remains close to that of $\mathbf{X}$, ensuring that key information is preserved.
3.7.3 Role of Nonlinear Activation in Enhancing Expressivity
While the Fourier transform is inherently linear, real-world data often exhibit nonlinear patterns. By using modReLU directly on the complex coefficients, we capture higher-order interactions that would otherwise necessitate more complex mechanisms in the token domain. The phase is preserved, and a threshold is applied to the magnitude, helping the model learn highly expressive frequency-based features.
@@ 3.8 Comparison with Convolution-Based Methods
It is important to emphasize that simply applying a global convolutional filter does not automatically equate to attention-like behavior. Although a global convolutional kernel can, in principle, span the entire sequence, it is typically a fixed or limited set of kernels shared across positions. In contrast, self-attention computes token-dependent (query-key) interactions that vary with the input sequence in a highly adaptive manner.
In our adaptive spectral filtering approach, the learnable filter $\mathbf{W}$ depends on the global context vector $\mathbf{c}$. Once $\mathbf{c}$ is extracted, the MLP-based modulation $\Delta\mathbf{W}$ alters the frequency filter in a data-dependent way, unlike most standard convolutional methods with fixed kernels. Furthermore, the $\operatorname{modReLU}$ nonlinear activation in the frequency domain can adaptively control which frequency components are kept or suppressed, effectively yielding different convolution kernels for different inputs. This leads to a form of dynamic convolution that is more akin to the adaptability seen in attention mechanisms, albeit realized entirely in the spectral domain.
In summary, while our method can be viewed through the lens of a global convolution (due to the convolution theorem relating frequency-domain multiplication to token-domain convolution), its key novelty lies in:
  • The global context-dependent filter $\mathbf{W}$.
  • The nonlinear modulation of complex coefficients via $\operatorname{modReLU}$.
  • The ability to adapt kernel shape according to the input.
These elements collectively endow our spectral filtering strategy with adaptive, attention-like capabilities that go beyond traditional convolutional architectures.
@@ 3.9 Comparison with FNet
FNet Lee-Thorp et al. (2022) also employs the DFT to mix tokens, removing all learnable parameters for mixing. However, it lacks adaptation to specific input distributions. Our method departs from that approach in key ways:
  • Adaptive Filtering: Instead of relying on a fixed DFT, we introduce a learnable filter conditioned on a global context vector, enabling input-dependent emphasis on certain frequencies.
  • Nonlinear Activation: We incorporate a complex-domain activation (modReLU) to capture higher-order phenomena that lie beyond the scope of linear transforms.
  • Strong Theoretical Basis: Our analysis uses energy preservation (Parseval’s theorem) and the convolution theorem to formally relate frequency-domain multiplication to global token mixing.
  • Practical Scalability: Although both methods achieve $\mathcal{O}(n\log n)$ complexity, our adaptive filtering and nonlinear activation introduce minimal overhead and substantially boost expressivity.
3.9.1 Computational Efficiency
Both FFT and IFFT run in $\mathcal{O}(n\log n)$ per channel. Our adaptive filtering and modReLU steps each add only $\mathcal{O}(n)$, so the overall cost remains dominated by the FFT and IFFT. This results in an efficient, scalable method for handling long sequences, standing in contrast to $\mathcal{O}(n^{2})$ self-attention mechanisms.
3.9.2 Mitigation Strategies for Adaptive MLP Complexity
The adaptive MLP in our spectral attention module generates modulation parameters for each frequency bin, where the number of bins is approximately $\frac{n}{2}+1$. Consequently, its output size scales linearly with the sequence length $n$, which may become a computational bottleneck for long sequences. We discuss several strategies to mitigate this issue:
  • Downsampling Frequency Bins: One effective approach is to reduce the number of frequency bins by applying pooling or subsampling techniques to the frequency-domain representation before feeding it into the adaptive MLP. This reduction preserves the essential spectral information while lowering the computational cost.
  • Parameter Sharing: Instead of generating independent modulation parameters for each frequency bin, the model can share parameters across groups of adjacent bins. This leverages the redundancy among nearby frequency components and significantly reduces the total number of parameters.
  • Low-Rank Approximations: By projecting the high-dimensional frequency representation into a lower-dimensional subspace using low-rank factorization, the adaptive MLP’s output size can be decreased. This strategy maintains the critical spectral features while reducing both the parameter count and computational overhead.
These techniques provide practical avenues to balance the trade-off between the expressive power of the adaptive MLP and its computational efficiency, especially when dealing with very long sequences.
In summary, adaptive spectral filtering:
  • Achieves global token mixing via the Fourier transform’s orthogonality.
  • Preserves signal energy for stable representations.
  • Approximates self-attention through a frequency-domain form of dynamic convolution.
  • Employs modReLU for higher-order interactions in the complex domain.
  • Operates in $\mathcal{O}(n\log n)$ time, offering a significant advantage over $\mathcal{O}(n^{2})$ attention-based methods.

### v4

@@ Abstract
Conventional self-attention mechanisms incur quadratic complexity, limiting their scalability on long sequences. We introduce FFTNet, an adaptive spectral filtering framework that leverages the Fast Fourier Transform (FFT) to achieve global token mixing in $\mathcal{O}(n\log n)$ time. By transforming inputs into the frequency domain, FFTNet exploits the orthogonality and energy preservation guaranteed by Parseval’s theorem to capture long-range dependencies efficiently. Our main theoretical contributions are 1) an adaptive spectral filter, 2) combining local windowing with a global FFT branch, and 3) rich nonlinearity introduction in both the frequency and token domains. Experiments on the Long Range Arena and ImageNet benchmarks validate our theoretical insights and demonstrate superior performance over fixed Fourier and standard attention models.
@@ 3 Adaptive Spectral Filtering Method
We present an adaptive spectral filtering framework that avoids explicit pairwise interactions while retaining global mixing capabilities. Rather than computing dot-product self-attention, this approach applies the discrete Fourier transform (DFT) to capture long-range dependencies in $\mathcal{O}(n\log n)$ time. By adaptively reweighting the resulting frequency components and then applying an inverse transform, the method achieves an effective balance of expressive power and computational efficiency, backed by solid theoretical foundations.
@@ 3.1 Motivation
Traditional attention mechanisms incur quadratic cost in the sequence length $n$ when computing pairwise token interactions. As $n$ increases, this quickly becomes expensive. By contrast, the DFT naturally encodes global interactions in $\mathcal{O}(n\log n)$ time, since it decomposes the token sequence into orthogonal frequency components. To increase representational power, we introduce a learnable filter in the frequency domain that emphasizes or attenuates particular frequency bands based on a global context.
Why Local Windowing?
While the global DFT branch is excellent for capturing large-scale dependencies, it can blur fine-grained or short-range signals. Local windowing (STFT) is optional, but it can help preserve the high-resolution structure within shorter segments. When used, merging these local and global representations yields a model capable of both wide-ranging context and precise local detail. If purely global context suffices, one may omit local windowing altogether.
@@ 3.2 Method Description
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be an input sequence of length $n$ and embedding dimension $d$. Our method integrates both global and local spectral representations to capture wide-ranging dependencies and fine-grained context. The approach proceeds in five main steps:
1. Global Fourier Transform.
Apply the discrete Fourier transform (FFT) across the entire sequence:
$\mathbf{F}_{\mathrm{global}}\;=\;\operatorname{FFT}(\mathbf{X})\;\in\;\mathbb{C}^{n\times d},$
which captures long-range dependencies without explicit pairwise comparisons.
2. Local Windowing (STFT). (Optional)
In parallel, and optionally, perform a short-time Fourier transform (STFT) for local context:
  1. 1.
    Partition the input $\mathbf{X}$ into overlapping windows of length $w$. Let each window be denoted by $\mathbf{X}^{(m)}\in\mathbb{R}^{w\times d}$, where $m$ indexes the window.
  2. 2.
    Apply a window function (e.g., Hann or Hamming) to each segment.
  3. 3.
    Compute the FFT for each window to obtain local frequency components:
    $\mathbf{F}_{\mathrm{local}}^{(m)}=\operatorname{FFT}\bigl{(}\mathbf{X}^{(m)}\bigr{)}.$
These local spectra emphasize short-range patterns that might be blurred in a purely global transform. If the application does not require high-resolution local detail, this step can be omitted.
3. Merging Global and Local Representations.
To maximize information flow from both branches, we propose an isometric fusion approach in the frequency domain. First, concatenate the global and local spectra along their feature dimension: [^1] If multiple local windows exist, they may be pooled or aggregated (e.g. average or max) before concatenation, or one can apply a deeper fusion network to combine them.
$\hat{\mathbf{F}}\;=\;\bigl{[}\mathbf{F}_{\mathrm{global}}\;\;\|\;\;\mathbf{F}_{\mathrm{local}}\bigr{]}\;\in\;\mathbb{C}^{n\times(2d)},$
where $\mathbf{F}_{\mathrm{local}}$ can be a representative fusion of all local windows. [^2] If retaining every window separately, this dimension becomes $n\times(d+\sum_{m}d_{\mathrm{local}}^{(m)})$.
Next, apply a learnable matrix $\mathbf{U}\in\mathbb{C}^{(2d)\times d}$ that fuses the concatenated features into a unified spectral representation:
$\mathbf{F}\;=\;\hat{\mathbf{F}}\,\mathbf{U}.$
In the simplest form, $\mathbf{U}$ can be unconstrained and learned end-to-end. However, to preserve (and thus maximize) the signal energy from both branches, $\mathbf{U}$ may be constrained to be isometric (or approximately isometric), satisfying $\mathbf{U}^{*}\mathbf{U}=\mathbf{I}$. This ensures that no information is lost during the fusion step, effectively providing a richer merge than a simple gating or element-wise sum.
@@
Compute a global context vector
$\mathbf{c}\;=\;\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i}$
and feed it into a small MLP to generate scaling and bias parameters, $\Delta\mathbf{s}$ and $\Delta\mathbf{b}$, for each frequency bin and attention head. These parameters modify a base filter and bias via
$\mathbf{W}\;=\;\mathbf{W}_{\mathrm{base}}\odot(1+\Delta\mathbf{s}),\quad\mathbf{b}\;=\;\mathbf{b}_{\mathrm{base}}+\Delta\mathbf{b},$
reweighting the complex coefficients:
$\tilde{\mathbf{F}}\;=\;\mathbf{F}\odot\mathbf{W}+\mathbf{b}.$
Here, $\mathbf{F}$ includes both local and global frequency information from the isometric fusion step.
5. Nonlinear Activation (modReLU) and Inverse FFT.
To enrich the representation, we apply modReLU Arjovsky et al. (2016) on each complex coefficient $z=re^{i\theta}$. This preserves the phase while applying a threshold to the magnitude:
$\operatorname{modReLU}(z)\;=\;\begin{cases}(r+b)\,e^{i\theta},&\text{if }r+b>0,\\[3.0pt] 0,&\text{otherwise}.\end{cases}$
Finally, we map back to the token domain via the inverse FFT:
@@
$\mathbf{Y}\;=\;\operatorname{IFFT}\bigl{(}\tilde{\mathbf{F}}\bigr{)}\;\in\;\mathbb{R}^{n\times d},$
yielding a globally mixed representation that also incorporates local spectral features (if local windowing was used).
3.3 Efficient Mixing
Using the FFT to encode global interactions has several advantages. It performs global mixing in $\mathcal{O}(n\log n)$ time, contrasting with the $\mathcal{O}(n^{2})$ cost of self-attention. Meanwhile, local STFT windows capture short-range structure that may be lost in a purely global transform. By isometrically merging local and global frequency components, we fuse high-resolution local context with a global view, all in roughly $\mathcal{O}(n\log n)$ time.
The adaptive filter serves as an implicit frequency-domain mask, conditioned on global context rather than explicit pairwise comparisons. Nonlinear activation in the complex domain (modReLU) further enriches representational capacity by capturing higher-order structure. Parseval’s theorem and the unitary properties of the DFT ensure that this transformation preserves both energy and inner-product structure.
3.4 Parseval’s Theorem
Parseval’s theorem asserts that the total energy of a signal remains constant under the Fourier transform, aside from a constant scaling factor. For an input sequence $\mathbf{X}\in\mathbb{R}^{n\times d}$ and its Fourier transform $\mathbf{F}=\operatorname{FFT}(\mathbf{X})$, the theorem states:
$\|\mathbf{X}\|_{2}^{2}\;=\;\frac{1}{n}\|\mathbf{F}\|_{2}^{2}.$
This energy preservation is crucial for our method as it guarantees that the adaptive filtering and subsequent nonlinear operations do not inadvertently distort the input signal’s inherent information.
3.5 Orthogonal Decomposition and Unitary Property of the DFT
The DFT can be represented by a matrix $\mathbf{F}_{n}\in\mathbb{C}^{n\times n}$ with entries given by
$[\mathbf{F}_{n}]_{j,k}\;=\;e^{-2\pi i(j-1)(k-1)/n}.$
This matrix is unitary up to a scaling factor, satisfying:
$\mathbf{F}_{n}^{*}\mathbf{F}_{n}\;=\;n\,\mathbf{I},$
where $\mathbf{F}_{n}^{*}$ is the conjugate transpose and $\mathbf{I}$ is the identity matrix. This property implies that the transformation preserves inner products:
$\langle\mathbf{X}_{i},\mathbf{X}_{j}\rangle\;=\;\frac{1}{n}\,\langle\mathbf{F}_{i},\mathbf{F}_{j}\rangle.$
Thus, every token in the input contributes to every frequency component, enabling effective global mixing akin to self-attention but at a significantly reduced computational cost.
@@ 3.6 Proof of Computational Complexity
Global FFT and Inverse FFT.
For $\mathbf{X}\in\mathbb{R}^{n\times d}$, the full-sequence FFT and IFFT each cost $\mathcal{O}(n\log n)$ per channel, leading to $\mathcal{O}(d\,n\log n)$ overall.
Local Windowing (STFT). (Optional)
If each window has length $w$ and we assume an overlap of $\alpha w$ tokens, the number of windows is approximately $\frac{n}{(1-\alpha)w}$. Each windowed FFT thus costs $\mathcal{O}(w\log w)$ for each of the $d$ channels. The total cost is:
$\mathcal{O}\!\Bigl{(}\tfrac{n}{(1-\alpha)w}\;d\;w\log w\Bigr{)},$
which is typically smaller or comparable to $\mathcal{O}(d\,n\log n)$ for practical choices of $w$. This step is optional and can be skipped if local structure is not critical for the task at hand.
Isometric Fusion and Adaptive Filtering.
Merging local and global representations with an isometric (or unconstrained) matrix $\mathbf{U}$ is an $\mathcal{O}(n\,d^{2})$ step when viewed as a matrix multiplication in $\mathbb{C}^{n\times(2d)}$. This is generally modest compared to the $\mathcal{O}(d\,n\log n)$ term for large $n$. Adaptive filtering, which includes the MLP for context-based parameter generation, similarly contributes $\mathcal{O}(d\,n)$.
@@
Because $\mathcal{O}(d\,n\log n)$ typically dominates the cost of element-wise operations for large $n$, the total complexity remains $\mathcal{O}(d\,n\log n)$. For practical configurations, this effectively behaves as $\mathcal{O}(n\log n)$.
Figure 1: Overall Process of FFTNet with Local Windowing (optional) and Global FFT, followed by isometric fusion in the frequency domain.
@@ 3.7 Comparison with Convolution-Based Methods
Although a sufficiently large convolution kernel can span the entire sequence, it generally lacks the adaptability of self-attention unless the kernel itself is dynamically generated. Our approach bridges this gap by using a data-dependent, learnable filter in the frequency domain, guided by global context and applied through modReLU. Because frequency-domain multiplication corresponds to convolution in the token domain, the model can realize global, context-aware mixing more flexibly than standard convolutional methods. Local windowing further refines the representation without resorting to large kernel sizes.
@@ 3.8 Comparison with FNet
FNet Lee-Thorp et al. (2022) similarly uses the DFT for token mixing but employs no learnable parameters in the mixing step. In contrast, our method includes a context-dependent, adaptive filter and a nonlinear activation that both enhance expressivity and allow emphasis on specific frequency components. Further, we incorporate a local STFT pathway to capture fine-grained interactions. Despite these adaptive steps, the overall complexity remains $\mathcal{O}(n\log n)$, dominated by the FFT itself.
3.8.1 Mitigating Complexity in the Adaptive MLP
For very long sequences, generating frequency-bin parameters can still be demanding. Possible solutions include (1) subsampling or pooling the frequency representation before the MLP, (2) sharing parameters among adjacent frequency bins, and (3) applying low-rank factorization to reduce dimensionality. These techniques help maintain a reasonable balance between the expressive power of the adaptive filter and computational cost.
Conclusion. Adaptive spectral filtering integrates frequency-domain global mixing, data-driven filtering, and nonlinear activation into a unified framework. By adding an optional local windowing path (STFT) and merging it with the global FFT via an isometric fusion, the model captures both short-range and long-range interactions with $\mathcal{O}(n\log n)$ efficiency. Parseval’s theorem and the orthogonal properties of the DFT ensure stable transformations, enabling many benefits of self-attention at a reduced computational cost. Full implementation details are available at: https://github.com/jacobfa/fft.